OpenStudy (anonymous):

If a car travels 340 miles on 20 gallons of gas, how far can it travel on 25 gallons of gas?

OpenStudy (anonymous):

340 = 20 x = 25 x = 25*340/20

OpenStudy (anonymous):

(25/ 20 ) * 340

OpenStudy (anonymous):

x = 425

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Xavier worked 10 hours on Monday and 15 hours on Wednesday. His total pay was $280.00. What is his rate per hour?

OpenStudy (anonymous):

x*10 + x*15 = 280 25*x = 280 x = 280/25 x = $11.2